Which of the following forms a molecular solid? palladium ° C, graphite QUESTION 5 Calculate the freezing point of a solution of 30.0 g methyl salicylate, C7H602, dissolved in 800. g of benzene, C6H6. Kf for benzene is 5.10°C/m and the freezing point is 5.50°C for benzene 0 1.56°C O -1.56°C 3.93°C O 7.06 C QUESTION 6 How many grams of H2 gas are there in a 5.00-L cylinder at 4.00 2.16 9 O 0.542 g O 823 g O 1.08 g x 10 mm Hg and 23°C? QUESTION 7 A supercritical fluid refers to a substance O above both its critical temperature and its critical pressure at its triple point. that is in the liquid crystal state o with a viscosity of zero
Answer #1

1) C, grapfhite is a molecular solid as well as covalent solid

2) Freezing point of solution:

\DeltaTF = KFM

No. of moles of methyl salicylate = wt taken/mol wt = 30/122= 0.246

Molality = No. of moles /wt of solvent in kg = 0.246/0.8 = 0.307

\DeltaTF = 0.307 x 5.10 =1.568 = Tf solvent - Tfsolution

-Tfsolution = \DeltaTF - Tf solvent = 1.568-5.50 = -3.93 C

Tfsolution = 3.93 C

3) PV = nRT 4000 mm of Hg = 5.26 atm T = 23+273 = 296 K

5X5.26 = nX 0.0820x 296

n = 26.3/24.30 = 1.082 = wt taken/mol wt

0.924x2= mass of hydrogen = 2.16 gms

4) A supercritical fluid refers to a substance above both its critica temperature and critical pressure.